# A Survey of Cognitive Distortion Detection and Classification in NLP

### v1

@@ 8 Limitations
As with any literature survey, our analysis is constrained by the scope, reporting quality, and coverage of the included studies, and should be viewed as a snapshot of a rapidly evolving field. As such, some very recent preprints may not be included. While we have aimed for balanced representation, our synthesis reflects our methodological choices and interpretive framing, which may influence the emphasis placed on particular themes. To support transparency and future research, we plan to release our paper list.
@@ 9 Ethical Considerations
Given the nature of cognitive distortions within psychotherapy contexts, this survey acknowledges several important ethical considerations. As our work is a meta-analysis and synthesis of existing studies, we did not collect new data or propose new models. Nevertheless, the scope of our review touches on certain areas of concern that warrant attention.
@@
For instance, with TherapistQA, the authors provide clear documentation - including detailed labelling guidelines and procedures for resolving disagreements between annotators (Shreevastava and Foltz, 2021). However, because the dataset draws on an upstream public source, some provenance details reflect the level of information made available by that original source rather than any omission by the curators themselves. In this case, the publicly available version is based on a Kaggle Q&A dataset [^3] https://www.kaggle.com/datasets/arnmaud/therapist-qa/data, for which we have not found publicly accessible details specifying the original platform or data collection process. This situation is not unique to TherapistQA; several widely used mental health corpora draw on similar repositories, highlighting a broader and ongoing challenge in achieving complete transparency of data origins within the field.
@@ Acknowledgments
Some cognitive distortion examples in Table 3 were generated with the assistance of ChatGPT. The tool was also used in its intended role as an academic support resource to refine and polish author-written text. All AI-assisted content was carefully reviewed, verified, and edited by the authors.
This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/W524475/1].
@@ Appendix A Survey Methodology
After manual filtering, we retained 38 primary publications and preprints, spanning from 2004 to May 2025. Studies were included if they (i) implemented a cognitive distortion detection or classification model, (ii) introduced a cognitive distortion-related dataset, or (iii) computationally explored the taxonomy of cognitive distortions. To ensure reproducibility, we will publicly release all our search queries, filtering criteria, and the resulting paper list online upon publication.
@@ B.2 Origins
The origins of cognitive distortions are often traced back to the early work of Beck in the 1960s. While he identified fewer types of distortions than are commonly recognised today, he did describe well-known types such as Overgeneralisation, as well as others that are less frequently cited. Quoted directly from his work (Beck, 1963) [^8] Both Arbitrary Interpretation and Selective Abstraction align reasonably well with the consolidated taxonomy reference in Table 1, corresponding to Jumping to Conclusions and Mental Filter, respectively.:
  • Arbitrary Interpretation – ‘the process of forming an interpretation of a situation, event, or experience when there is no factual evidence to support the conclusion, or when the conclusion is contrary to the evidence.’
  • Selective Abstraction – ‘focusing on a detail taken out of context, ignoring other more salient features of the situation, and conceptualising the whole experience on the basis of this element.’

### v2

@@ 9 Limitations
As with any literature survey, our analysis is constrained by the scope, reporting quality, and coverage of the included studies, and should be viewed as a snapshot of a rapidly evolving field. As such, some very recent preprints may not be included. While we have aimed for balanced representation, our synthesis reflects our methodological choices and interpretive framing, which may influence the emphasis placed on particular themes.
@@ 10 Ethical Considerations
Given the nature of cognitive distortions within psychotherapy contexts, this survey acknowledges several important ethical considerations. As our work is a synthesis of existing studies, we did not collect new data or propose new models. Nevertheless, the scope of our review touches on certain areas of concern that warrant attention.
@@
For instance, with TherapistQA, the authors provide clear documentation - including detailed labelling guidelines and procedures for resolving disagreements between annotators (Shreevastava and Foltz, 2021). However, because the dataset draws on an upstream public source, some provenance details reflect the level of information made available by that original source rather than any omission by the curators themselves. In this case, the publicly available version is based on a Kaggle Q&A dataset, [^6] https://www.kaggle.com/datasets/arnmaud/therapist-qa/data for which we have not found publicly accessible details specifying the original platform or data collection process. This situation is not unique to TherapistQA; several widely used mental health corpora draw on similar repositories, highlighting a broader and ongoing challenge in achieving complete transparency of data origins within the field.
@@ Acknowledgments
This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/W524475/1]. We thank the anonymous reviewers for their constructive feedback.
@@ Appendix A Survey Methodology
After manual filtering, we retained 38 primary publications and preprints, spanning from 2004 to May 2025. Studies were included if they (i) implemented a cognitive distortion detection or classification model, (ii) introduced a cognitive distortion-related dataset, or (iii) computationally explored the taxonomy of cognitive distortions. To support reproducibility, we release our supplementary resources and paper list, along with corrections and updates, on GitHub: https://github.com/archiesage/cognitive-distortion-nlp-survey.
@@ B.2 Origins
The origins of cognitive distortions are often traced back to the early work of Beck in the 1960s. While he identified fewer types of distortions than are commonly recognised today, he did describe well-known types such as Overgeneralisation, as well as others that are less frequently cited. Quoted directly from his work (Beck, 1963): [^11] Both Arbitrary Interpretation and Selective Abstraction align reasonably well with the consolidated taxonomy reference in Table 1, corresponding to Jumping to Conclusions and Mental Filter, respectively.
  • Arbitrary Interpretation - ‘the process of forming an interpretation of a situation, event, or experience when there is no factual evidence to support the conclusion, or when the conclusion is contrary to the evidence.’
  • Selective Abstraction - ‘focusing on a detail taken out of context, ignoring other more salient features of the situation, and conceptualising the whole experience on the basis of this element.’